radomes.openskiesproject.org
AdoreDior AdoreDior
  • 02-12-2015
  • Mathematics
Answered

A fish tank contains 18 goldfish and 22 guppies. If you randomly select 2 fish, what is the probability that they are both goldfish?

Verified Answer :

redcloud redcloud
  • 02-12-2015

18 goldfish+22 guppies=40 total
18/40 are goldfish and 22/40 are guppies
Odds that 1 goldfish is chosen at random is 18/40 or 9/20 when simplified so odds that 2 goldfish are chosen at random is 9/20×9/20=81/400 or 20.25%

Answer Link

Related Questions

How far west did the Muslim Empire spread at its height
speeding peeling out or weaving through traffic are examples of
How did Congress approach Reconstruction after the Civil War
Which of Americas characteristics led England to colonize there Areligious intolerance and high rents Babundant resources and an opportunity for land ownership
Given a 3 b 4 and c 5 evaluate a b c
is it possible for water to become stuck in a step of the water cycle
each interior angle of a regular polygon measures 156 how many sides does the polygon have
The number of cars c in a parking lot increases when the parking fee f decreases Write the correct equation for this scenario and solve for the number of cars w
Which accurately describes the lines in each stanza of the poem I Am Offering This Poem
speeches contain rhyme and a meter that is different from the speech of other characters a Lady Macbeths b Banquos c The murderers d Hecates